do pair of staight lines always intersect?if no then how do we know whether parallel pair of straight lines are homogeneous or not

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Analgebric eqn. is homogenous eqn. if the sum of the powers of x and y in each term is equal to 2.
 
ax^2 + 2hxy + by^2 = 0
is the general form of a homogeneous eqn. of second degree in x and y and it represents strt. lines passing through the origin and these lines will be real and distinct if 
  1. h^2 > ab
  2. real aand coincide if h = ab
  3. if h^2 = ab then these lines will be coincide.
  4. if a+b =0 then two lines will be perpendicular to each other